OpenStudy (anonymous):

A channel is to be constructed out of a sheet of metal, to allow gravel to be removed from a mining site. The sheet of metal has a width 40cm, and each side is to be bent up (at the same angle theta=pi/4 from the horizontal) so as to maximise the amount of gravel that the channel can carry

OpenStudy (anonymous):

Have been working on it all day still got no-where.

OpenStudy (whpalmer4):

Can you write an expression for the cross-sectional area as a function of \(w\)?

OpenStudy (anonymous):

Have to find the portion -w- on the sheet metal as well as calculate maximum cross sectional area of the channel.

OpenStudy (anonymous):

The length of the top side is: (2wcos(45) + L) = 2 * w / sqrt(2) + L Area of trapezium A = 1/2 * (L + 2*w/sqrt(2) + L) * w / sqrt(2) A = 1/2 * (2L + 2*w/sqrt(2)) * w/sqrt(2) = (L + w/sqrt(2)) * w /sqrt(2) A = Lw/sqrt(2) + w^2/2 Put L = 40 - 2w in A. A = w/sqrt(2) * (40 - 2w) + w^2/2 THIS is as far as ive gotten.

OpenStudy (whpalmer4):

Okay, I agree with that expression of the area. Do you recognize the form of the equation?

OpenStudy (whpalmer4):

It's a parabola. Find the vertex and you've got your solution. Vertex of a parabola in the form \(ax^2+bx+c\) is at \(x = -\dfrac{b}{2a}\)

OpenStudy (anonymous):

The vertex gives me my maximum cross sectional area correct?

OpenStudy (whpalmer4):

with the equation written as you have it, the vertex of the parabola is at the value of \(w\) that gives the maximum cross-sectional area

OpenStudy (anonymous):

Slightly confused. How do i isolate the two?

OpenStudy (whpalmer4):

isolate which two?

OpenStudy (anonymous):

Ive been told to continue with that, find the derivative and make it equate to 0 to get max. Then solve for w.

OpenStudy (whpalmer4):

well, okay, if you want to do it the longer way, but for any parabola of the form \(y = ax^2+bx+c\), the vertex is at \(x = -\frac{b}{2a}\) which should be able to convince yourself of by taking the derivative of the parabola equation, setting it equal to 0 and solving for \(x\).

OpenStudy (anonymous):

Can you show me how id go by doing that? still confused. sorry.

OpenStudy (whpalmer4):

You have \[A = \frac{1}{2}w^2-\sqrt{2}w^2 + 20\sqrt{2} w = (\frac{1}{2}-\sqrt{2})w^2 + 20\sqrt{2}w\]\[a = (\frac{1}{2}-\sqrt{2})\]\[b = 20\sqrt{2}\]vertex is located at \[w=-\frac{20\sqrt{2}}{2(\frac{1}{2}-\sqrt{2})}\]sharpen your pencil and simplify that that...

OpenStudy (anonymous):

Giving w = 15.469 (3dp). Do i take this sub it back into my original formula of A = And allgood?

OpenStudy (whpalmer4):

yes, that would give you the cross-sectional area

OpenStudy (anonymous):

1407.827 sound correct?

OpenStudy (whpalmer4):

Here's a plot of the area function as w goes from 0 to 31. No, 1408 does not sound correct. If we had a rectangular channel of width 40, it would have to be approximately 35 tall to have that cross-sectional area!

OpenStudy (anonymous):

106.857(3dp)?

OpenStudy (whpalmer4):

no. how are you evaluating this?

OpenStudy (whpalmer4):

Take \(w= 15.469\). Square it. Multiply it by \(a = \frac{1}{2}-\sqrt{2} \approx -0.914\) Take \(w = 15.469\). Multiply it by \(b = 20\sqrt{2} \approx 28.284\) Add the two together. What do you get?

OpenStudy (whpalmer4):

compare your answer with the y-value of the vertex of the parabola in the graph I gave you. the vertical line indicates the position of the vertex, if it isn't clear

OpenStudy (whpalmer4):

it's like 4AM here. time to hit the sack. good luck wrapping up the problem!

OpenStudy (anonymous):

209.012

OpenStudy (anonymous):

^ WOW ahaha.

OpenStudy (whpalmer4):

still off by about 10, I'm not sure what you're doing wrong

OpenStudy (anonymous):

No worries. Thanks for your help.